# Cover Learning for Large-Scale Topology Representation

### v1

@@ 4.4 ShapeDiscover
See Algorithm 3 for a high-level description of the procedure; here we give details about the subroutines. For details about the software we use, see Section G.1.
$\mathrm{NeighborhoodGraph}$. The simplest approach is to use a $k$-nearest neighbor graph, with weights equal to $1$. This worked well in most of our experiments. However, to make the method more robust, our implementation uses the weighted neighborhood graph of the UMAP algorithm (Section 3.1 (McInnes et al., 2018)), and this is what we use for the computational examples.
@@
  Input: Point cloud $X\subseteq\mathbb{R}^{N}$
[…]
  Optimization parameters: lr, n_epoch, $p\in[1,\infty)$
[…]
  $\theta^{\prime}\leftarrow\mathrm{InitializeParametricModel}\left(h_{(-)},g\right)$
@@
  $\theta\leftarrow\mathrm{GradientDescent}(\mathcal{L},\texttt{n\_epoch},\texttt{lr},init=\theta^{\prime})$
  Return $\pi_{\infty}\circ h_{\theta}$
[…]
$\mathrm{InitializeParametricModel}$. Since we use a very simple model, initialization is straightforward, and it amounts to setting $\theta^{\prime}_{i,j}\leftarrow g(x_{i})_{j}$.
$\mathrm{GradientDescent}$. We train for n_epoch (or until convergence) using an Adam optimizer with learning rate lr and otherwise default parameters.
@@ 4.5 Output and parameter selection
We elaborate on how to use ShapeDiscover’s output, and how we choose its parameters in the experiments.
Output. The output of Algorithm 3 is a fuzzy cover. Using the nerve construction (Definition A.1), the fuzzy cover induces a filtered simplicial complex (Appendix C), indexed by $[0,1]$. Given any $\lambda\in[0,1]$, the filtered simplicial complex can be thresholded to a single simplicial complex.
[…]
When a filtered simplicial complex is needed, we reindex the output of Algorithm 3 by $-\log$ to have it be indexed by $[0,\infty)$ as standard topological inference algorithms (Appendix C). When a simplicial complex is needed, we threshold with $\lambda=0.5\in[0,1]$, by default.
@@ 5 Experiments
See appendix for details about software (Section G.1), datasets (Section G.2), and computation times (Table 4).
[…]
@@ 5.1 Topological inference
A. Efficient topological representation. We compare ShapeDiscover to the standard approaches to topological inference on four datasets: a synthetic 2-sphere, a synthetic 3-sphere, a sample of the surface of a human (topological 2-sphere), and the dynamical system video data from (Lederman & Talmon, 2018) (topological 2-torus). We do not use Mapper since current implementations are not suitable for higher-dimensional topological inference (Section G.5). Each algorithm takes as input a number of vertices, and the goal is to find the smallest number of vertices that achieves a certain homology recovery quotient, quantified as the proportion of filtration values at which the correct homology is recovered (see Section G.4 for the definition). For more details; see Section G.3. Results are in Tables 1 and 2. ShapeDiscover performs best, followed by Witness ($v=1$).
B. Toroidal topology in neuroscience. In Section G.6, we describe an experiment on the toroidal topology of population activity in grid cells in the dataset of (Gardner et al., 2022). In it, ShapeDiscover is able to learn this topology using a simplicial complex with just $15$ vertices (see Figure 6), a task that Vietoris–Rips is not able to accomplish with a simplicial complex with $5000$ vertices.
@@ 5.2 Visualization of large-scale structure
For plotting conventions, see Section G.7.
C. Human and octopus. We compare ShapeDiscover to 1D Mapper (height as filter function), Ball Mapper, and Differentiable Mapper on human and octopus datasets (Figure 7). In Figs. 1,2 ShapeDiscover recovers the large-scale topology, including the fact that the figures are hollow, with a small number of vertices, while 1D Mapper and Differentiable Mapper recover the $1$-dimensional structure, but not the higher-dimensional one (Section 2). In Figures 11 and 12 Ball Mapper has a main parameter ($\varepsilon$) that is hard to tune, and has a hard time recovering the correct topology.
@@
D. Handwritten digit datasets. We run ShapeDiscover and UMAP on a small handwritten digits dataset and on MNIST. In Figures 4 and 15 UMAP and ShapeDiscover capture similar structure, corresponding well to digits and their similarities. We did not obtain meaningful results with other algorithms (Figures 14 and G.8).
E. Flare structure in biology. We use the data from (Packer et al., 2019), consisting of the gene expressions of cells of the C. elegans worm, and known to have flare structure corresponding to single-cell trajectories. We project the data to 2 dimensions using UMAP, which clearly identifies the flare structure of $16$ cell types (111116161616); see Figures 16 and 3. ShapeDiscover identifies the same flare structure (Figure 5). We were not able to obtain meaningful results with other algorithms (Figures 17 and G.8).
@@ 6 Conclusions
Motivated by the problem of representing the large-scale topology of geometric data for the purposes of efficient topological inference and visualization, we have developed a theory for cover learning based on optimization, and an algorithm which outperforms state-of-the-art methods in TDA for both purposes, with easy-to-tune parameters.
We believe that cover learning might be interest to the wider unsupervised learning community, with the potential of having applications in clustering and dimensionality reduction.
This is only a first step towards the development of efficient and robust cover learning methods, opening up several research directions; we mention a few below.
@@
Limitations and future work. Covers produced by cover learning algorithms will often contain spurious intersections, which can render visualization difficult, and lead to incorrect topology recovery. Robust cover learning, which incorporates density structure, and the development of reliable and interpretable layout algorithms for the nerve of covers are left to future work.
@@
Standard topological inference algorithms, such as Vietoris–Rips, come with topology recovery convergence guarantees (e.g., (Kim et al., 2020)), and here we have not provided such guarantees for optimization-based cover learning procedures. We believe that the nerve theorem can be used to state computable conditions which, when satisfied, guarantee topology recovery of covers, regardless of how they were obtained.
@@
Topological persistence optimization is slow for various reasons, including the fact that the whole dataset is required at each iteration step, precluding the usage of mini-batch for more efficient optimization. It is an open question whether there exist methods for mini-batch topological regularization (i.e., enforcing trivial topology, which is what the nerve theorem requires). It may be possible to develop such methods using probabilistic algorithms for topological estimation.
@@
Table 1: Rows correspond to algorithms taking a number of vertices and outputting a filtered complex. Columns correspond to a dataset and a homology recovery quotient. Entries are the minimum number of vertices (left) and total number of simplices (right) needed to get the required homology recovery quotient; see Section G.3 for further details.
@@
[…]
@@
[…]
@@
[…]
@@ A.1 Simplicial complexes and nerve
A finite simplicial complex consists of a finite set $K_{0}$ of vertices, together with a set $K$ of subsets of $K_{0}$ with the property that $\{x\}\in K$ for every $x\in K_{0}$, and such that, for all $\sigma\subseteq\tau\in K$, we have $\sigma\in K$. The elements of $K$ are called simplices, the dimension of a simplex $\sigma\in K$ is $|\sigma|-1\in\mathbb{N}$, and the dimension of a simplicial complex is the maximum dimension of its simplices. For example, any finite, unoriented, simple graph $G$ can equivalently be seen as a $1$-dimensional simplicial complex with 00-dimensional simplices the vertices of $G$, with $1$-dimensional simplices the edges of $G$, and with no higher-dimensional simplices.
@@ A.2 Homology, homology-good covers, and nerve theorem
We fix a field, which we omit from notation. The homology construction takes as input a topological space $X$, and returns vector spaces $\mathsf{H}_{i}(X)$ for each $i\in\mathbb{N}$, in such a way that $\dim(\mathsf{H}_{i}(X))$ counts the number of $i$-dimensional holes in $X$; for example, $\mathsf{H}_{0}(X)$ counts connected components, while $\mathsf{H}_{1}(X)$ counts essential loops. When referring to all homology groups at once, we write $\mathsf{H}(X)$.
@@
Let $\mathcal{U}=\{U_{i}\}_{i\in I}$ be an open cover of a topological space $X$, and let $k\in\mathbb{N}$. The cover $\mathcal{U}$ is $\ell$-homology-good if $\overline{\mathsf{H}}_{m}\left(\bigcap_{j\in J}U_{j}\right)=0$ for every $J\subseteq I$ with $|J|\leq\ell+1$ and $10\leq m\leq\ell-|J|-$ 1.
@@ A.3.2 Other indexing posets
We remark that geometric complexes used in topological inference, such as the Vietoris–Rips complex (Section A.4) are often indexed by the non-negative real numbers. Another indexing poset of interest is $([0,1],\geq)$, which indexes the filtered simplicial complexes associated to fuzzy covers; see Appendix C.
@@ A.5 Topological persistence optimization
There is extensive literature on the optimization of losses based on topology (Carrière et al., 2021; Leygonie et al., 2023, 2022; Hensel et al., 2021), with applications such as classification (Hofer et al., 2017; Poulenard et al., 2018; Hofer et al., 2020b; Horn et al., 2022; Zhao et al., 2020) representation learning (Hofer et al., 2019), and regularization (Chen et al., 2019; Hofer et al., 2020a; Scoccola et al., 2024). Of particular importance to this paper is the idea of topological optimization with big steps introduced in (Nigmetov & Morozov, 2024), which addresses the sparsity of gradient in vanilla topological persistence optimization, making gradient descent convergence significantly faster than previous approaches (for another, recent approach see (Carrière et al., 2024)). For more details, see Section A.3.
@@ B.1 Computation of the loss function
[…]
@@ B.2 Relaxation of the topological loss
[…]
@@ B.3 The probability measure associated to a fuzzy cover
There is theory of random sets, which deals with this type of situation; see, e.g., (Molchanov, 2017) or Chapter 18 (Aliprantis & Border, 2006). Here, we shall not go into details, since the functionals of the random covers we are interested in can be expressed as explicit integrals which do not involve random sets; see Theorem 3.2.
Appendix C The filtered simplicial complex associated to a fuzzy cover

### v2

@@ 4.5 ShapeDiscover
See Algorithm 3 for a high-level description of the procedure; here we give details about the subroutines. For details about the software we use see Section I.1, and for computational complexity see Appendix C.
Subroutine $\mathrm{NeighborhoodGraph}$
The simplest approach is to use a nearest neighbor graph, with weights equal to $1$. This worked well in most of our experiments. However, to make the method more robust, our implementation uses the weighted neighborhood graph of the UMAP algorithm (see Section 3.1 of (McInnes et al., 2018)); this is what we use in the computational examples.
@@
Input: Point cloud $X\subseteq\mathbb{R}^{N}$.
Parameters: $\texttt{n\_cov}\in\mathbb{N}$, $\texttt{n\_neigh}\in\mathbb{N}$, $\texttt{reg}>0$.
Optimization parameters: lr, n_epoch, $p\in[1,\infty)$.
$G:=\mathrm{NeighborhoodGraph}(X,\texttt{n\_neigh})$
$g:=\mathrm{FuzzyCoverInitialization}(G,\texttt{n\_cov})$
$h:=\mathrm{ParametricPartitionOfUnity()}$
$\theta^{\prime}:=\mathrm{InitializeParametricModel}\left(h,g\right)$
@@
$\theta:=\mathrm{GradientDescent}(\mathcal{L},\texttt{n\_epoch},\texttt{lr},init=\theta^{\prime})$
Return $\pi_{\infty}\circ h_{\theta}$.
Subroutine $\mathrm{FuzzyCoverInitialization}$
We use spectral clustering as initialization: We compute the first n_cov eigenvalues of the normalized Laplacian of $G$, and apply $k$-means with $k=\texttt{n\_cov}$ to get a clustering of $X$, which we then simply interpret as a fuzzy cover, as described in Section 4.4.
Subroutine $\mathrm{ParametricPartitionOfUnity}$
We use the simplest of the approaches described in Section 4.2: we let $\varphi$ be a $|X|$-by-n_cov real matrix, and let $f_{\theta}:X\to\mathbb{R}^{k}$ be defined by $f_{\theta}(x_{i})_{j}=\theta_{i,j}$. Then, the parametric partition of unity is defined as $\mathsf{softmax}\circ f_{\theta}:X\to\Delta^{k-1}_{1}$.
Subroutine $\mathrm{InitializeParametricModel}$
Since we use a very simple model, initialization is straightforward, and it amounts to setting $\theta^{\prime}_{i,j}\leftarrow g(x_{i})_{j}$.
Subroutine $\mathrm{GradientDescent}$
We train for n_epoch (or convergence) using an Adam optimizer with learning rate lr and default parameters.
@@ 4.6 Output and parameter selection
We elaborate on how to use ShapeDiscover’s output, and on how we choose its parameters in the experiments. See also Section I.9 for ablation study and parameter and noise sensitivity analyses.
4.6.1 Output
The output of Algorithm 3 is a fuzzy cover. Using the nerve construction (Definition A.1), the fuzzy cover induces a filtered simplicial complex (Appendix D), indexed by $[0,1]$. Given any $\lambda\in[0,1]$, the filtered simplicial complex can be thresholded to a single simplicial complex.
4.6.2 Default parameters
In all experiments we fix the following default parameters: number of neighbors for the neighborhood graph $\texttt{n\_neigh}=15$, regularization parameter $\texttt{reg}=10$, number of iterations for gradient descent $\texttt{n\_epoch}=500$, learning rate for gradient descent $\texttt{lr}=0.1$, and approximation parameter for fuzzy cover $p=5$ (Section 4.4).
When a filtered simplicial complex is needed, we reindex the output of Algorithm 3 by $\log$ to have it be indexed by $0,\infty)$ as standard topological inference algorithms (Appendix D). When a simplicial complex is needed, we threshold with $\lambda=0.5\in[0,1]$, by default.
@@ 5 Experiments
See appendix for details about software (Section I.1), datasets (Section I.2), computation time (Table 4), computational complexity (Appendix C), as well as an ablation study and parameter and noise sensitivity analyses (Section I.9).
We use ShapeDiscover with default parameters in all experiments (Section 4.6). We only vary the maximum cover size ($\texttt{n\_cov}\in\mathbb{N}$) and threshold $\lambda\in[0,1]$ (but only in MNIST and C. elegans data, otherwise default $\lambda=0.5$ is used).
@@ 5.1 Topological inference
A. Efficient topological representation
We compare ShapeDiscover to the standard approaches to topological inference on four datasets: a synthetic 2-sphere, a synthetic 3-sphere, a sample of the surface of a human (which is topologically equivalent to a 2-sphere), and the dynamical system video data from (Lederman & Talmon, 2018) (which is topologically equivalent to a 2-torus). We do not use Mapper since current implementations are not suitable for higher-dimensional topological inference (Section I.5). Each algorithm takes as input a number of vertices, and the goal is to find the smallest number of vertices that achieves a certain homology recovery quotient, quantified as the proportion of filtration values at which the correct homology is recovered (see Section I.4 for the definition). For more details; see Section I.3. Results are in Tables 1 and 2. ShapeDiscover performs best, followed by Witness ($v=1$).
B. Toroidal topology in neuroscience
In Section I.6, we describe an experiment on the toroidal topology of population activity in grid cells in the dataset of (Gardner et al., 2022). In it, ShapeDiscover is able to learn this topology using a simplicial complex with just $15$ vertices (see Figure 6), a task that Vietoris–Rips is not able to accomplish with a simplicial complex with $5000$ vertices.
@@ 5.2 Visualization of large-scale structure
For plotting conventions, see Section I.7.
C. Human and octopus
We compare ShapeDiscover to 1D Mapper (height as filter function), Ball Mapper, and Differentiable Mapper on human and octopus datasets (Figure 7). In Figures 1 and 2, ShapeDiscover recovers the large-scale topology, including the fact that the figures are hollow, with a small number of vertices, while 1D Mapper and Differentiable Mapper recover the $1$-dimensional structure, but not the higher-dimensional one (Section 2). In Figures 11 and 12, Ball Mapper has a main parameter ($\varepsilon$) that is hard to tune, and has a hard time recovering the correct topology.
@@ C. Human and octopus
D. Handwritten digit datasets
We run ShapeDiscover and UMAP on a small handwritten digits dataset and on MNIST. In Figures 4 and 15 UMAP and ShapeDiscover capture similar structure, corresponding well to digits and their similarities. We did not obtain meaningful results with other algorithms (Figures 14 and I.8).
E. Flare structure in biology
We use the data from (Packer et al., 2019), consisting of the gene expressions of cells of the C. elegans worm, and known to have flare structure corresponding to single-cell trajectories. We project the data to 2 dimensions using UMAP, which clearly identifies the flare structure of $16$ cell types (111616); see Figures 16 and 3. ShapeDiscover identifies the same flare structure (Figure 5). We were not able to obtain meaningful results with other algorithms (Figures 17 and I.8).
@@ 6 Conclusions
Motivated by the problem of representing the large-scale topology of geometric data for the purposes of efficient topological inference and visualization, we have developed a theory for cover learning based on optimization, as well as a practical implementation of these ideas with easy-to-tune parameters. We demonstrated that cover learning can improve upon the two main TDA methodologies with a single approach, by effectively addressing their main shortcomings: the sometimes prohibitive size of geometric complexes, and the difficulty in tuning and the lack of higher dimensional information in Mapper graphs.
We believe that cover learning is relevant to unsupervised learning, and to clustering and dimensionality reduction specifically. This paper is only a first step towards the development of efficient and robust cover learning methods. We now elaborate on the limitations of our approach, and give further motivation for future research directions.
@@
Robust cover learning. Covers produced by cover learning algorithms will often contain spurious intersections, which can render visualization difficult, and lead to incorrect topology recovery. This motivates the developing of robust cover learning, and of reliable and interpretable layout algorithms for cover visualization.
@@
Consistency. Standard topological inference algorithms come with topology recovery guarantees, e.g., (Boissonnat et al., 2018; Kim et al., 2020), and here we have not provided such guarantees for optimization-based cover learning procedures. As an avenue for future work, we observe that the nerve theorem can be used to formulate computable conditions for covers which, when satisfied, guarantee topological correctness of the cover, regardless of how it was obtained.
@@
Efficiency. Topological persistence optimization is slow for a few reasons, including the fact that the whole dataset is required at each iteration step, precluding the usage of mini-batch during optimization. It is an open problem to find methods for topological regularization (i.e., for enforcing trivial topology), which can be computed in small batches of the data. It may be possible to develop such methods using probabilistic, sample-based algorithms for topological estimation.
Other cover learning algorithms. Cover learning being a general unsupervised learning problem (like clustering), there is no single, ultimate cover learning algorithm, and other approaches should be considered. In particular, one could simplify our pipeline, as suggested by the ablation study (Section I.9).
Point cloud vectorization. Point clouds can be vectorized by constructing a graph on the data and applying methods from graph machine learning. Cover learning can be used to build graphs on data that are of much smaller size than those built using classical methods such as $k$-nearest neighbors, effectively reducing the dimension of the learning problem.
Persistence-based clustering. A main difficulty with many clustering algorithms is that the final number of clusters needs to be chosen by the user. This difficulty can be addressed with methods from density-based clustering and persistence-based clustering (Campello et al., 2013; McInnes et al., 2017). Fuzzy cover learning naturally produces a filtered graph (Section 4.6.1), which can then be used as input to persistence-based clustering methods such as (Chazal et al., 2013; Rolle & Scoccola, 2024).
@@
Table 1: Rows correspond to algorithms taking a number of vertices and outputting a filtered complex. Columns correspond to a dataset and a homology recovery quotient. Entries are the minimum number of vertices (left) and total number of simplices (right) needed to get the required homology recovery quotient; see Section I.3 for further details.
@@
Figure 2: ShapeDiscover ($\texttt{n\_cov}=20$) (left) and 1D Mapper (right) on octopus dataset. See Figs. 8,9 for corresponding covers.
@@
Figure 3: ShapeDiscover ($\texttt{n\_cov}=15$) (left) and UMAP (right) on the small digits dataset; see Figure 13 for ShapeDiscover’s cover on the UMAP projection. We see that both algorithms capture essentially the same structure: most of the digits are clearly distinguished, except for 88 and $1$. Note that only $11$ of the $15$ elements of ShapeDiscover’s cover are non-empty.
@@
Figure 4: ShapeDiscover ($\texttt{n\_cov}=15$, $\lambda=0.8$) (left) and UMAP (right) on MNIST; see Figure 15 for ShapeDiscover’s cover on the UMAP projection. The algorithms capture similar structure; a main difference is that ShapeDiscover isolates a large subset of the digit 11 (vertex I). Note that only $10$ of the $15$ elements of ShapeDiscover’s cover are non-empty.
Figure 5: ShapeDiscover ($\texttt{n\_cov}=100$, $\lambda=0.6$) on C. elegans data, recovering most of the flare structure captured by UMAP (Figures 16 and 3); $53$ elements in the cover are non-empty.
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgments
@@ A.1 Simplicial complexes and nerve
A finite simplicial complex consists of a finite set $K_{0}$ of vertices, together with a set $K$ of subsets of $K_{0}$ with the property that $\{x\}\in K$ for every $x\in K_{0}$, and such that, for all $\sigma\subseteq\tau\in K$, we have $\sigma\in K$. The elements of $K$ are called simplices, the dimension of a simplex $\sigma\in K$ is $|\sigma|-1\in\mathbb{N}$, and the dimension of a simplicial complex is the maximum dimension of its simplices. For example, any finite, unoriented, simple graph $G$ can equivalently be seen as a $1$-dimensional simplicial complex with 0-dimensional simplices the vertices of $G$, with $1$-dimensional simplices the edges of $G$, and with no higher-dimensional simplices.
@@ A.2 Homology, homology-good covers, and nerve theorem
We fix a field, which we omit from notation. The homology construction takes as input a topological space $X$, and returns a vector space $\mathsf{H}_{i}(X)$ for each $i\in\mathbb{N}$, in such a way that $\dim(\mathsf{H}_{i}(X))$ counts the number of $i$-dimensional holes in $X$; for example, $\mathsf{H}_{0}(X)$ counts connected components, while $\mathsf{H}_{1}(X)$ counts essential loops. When referring to all homology groups at once, we write $\mathsf{H}(X)$.
@@
Let $\mathcal{U}=\{U_{i}\}_{i\in I}$ be an open cover of a topological space $X$, and let $k\in\mathbb{N}$. The cover $\mathcal{U}$ is $\ell$-homology-good if $\overline{\mathsf{H}}_{m}\left(\bigcap_{j\in J}U_{j}\right)=0$ for every $J\subseteq I$ with $|J|\leq\ell+1$ and $0\leq m\leq\ell-|J|-1$.
@@ A.3.2 Other indexing posets
We remark that geometric complexes used in topological inference, such as the Vietoris–Rips complex (Section A.4) are often indexed by the non-negative real numbers. Another indexing poset of interest is $([0,1],\geq)$, which indexes the filtered simplicial complexes associated to fuzzy covers; see Appendix D.
@@ A.5 Topological persistence optimization
There is extensive literature on the optimization of losses based on topology (Carrière et al., 2021; Leygonie et al., 2023, 2022; Hensel et al., 2021), with applications such as classification (Hofer et al., 2017; Poulenard et al., 2018; Hofer et al., 2020b; Horn et al., 2022; Zhao et al., 2020) representation learning (Hofer et al., 2019), and regularization (Chen et al., 2019; Hofer et al., 2020a; Scoccola et al., 2024; Moor et al., 2020). Of particular importance to this paper is the idea of topological optimization with big steps introduced in (Nigmetov & Morozov, 2024), which addresses the sparsity of gradient in vanilla topological persistence optimization, making gradient descent convergence significantly faster than previous approaches (for another, recent approach see (Carrière et al., 2024)). For more details, see Section A.3.
@@ B.1 Computation of the loss function
For the inequality, note that $\mathbb{P}\left(\mathcal{U}\text{ is not homology-good}\right)=\mathbb{E}(\psi_{\mathcal{U}})$, where $\psi_{\mathcal{U}}$ is 0 is $\mathcal{U}$ is good, and $1$ otherwise. By the considerations in the first paragraph, it follows that
$\mathbb{P}\left(\mathcal{U}\text{ is not homology-good}\right)\;$ $=\;\int_{0}^{1}\psi_{\left\{g>\lambda\right\}}\,\mathsf{d}\lambda$
@@ B.2 Relaxation of the topological loss
Let $g:\mathcal{M}\to\Gamma^{k-1}$ be a fuzzy cover, and let $\mathcal{U}=\{U_{i}\}_{i=1}^{k}$ be a random cover distributed as $\mu_{g}$, as in 1. Let $I=\{1,\dots,k\}$. Fix a maximum homology dimension $\ell\in\mathbb{N}$ of interest. Then, the topological loss $\mathsf{T}(g)=\mathbb{P}\left(\left\{U_{i}\right\}\text{ is not homology-good }\right)$ can be relaxed to a loss $\mathsf{T}_{\ell}(g)=\mathbb{P}\left(\left\{U_{i}\right\}\text{ is not $\ell$-homology-good }\right)$, using Definition A.4. A derivation analogous to that in the proof of Theorem 3.2 shows that
@@ B.3 The probability measure associated to a fuzzy cover
There is a more general theory of random sets, which deals with this type of situation; see, e.g., (Molchanov, 2017) or Chapter 18 (Aliprantis & Border, 2006). Here, we shall not go into details, since the functionals of the random covers we are interested in can be expressed as explicit integrals which do not involve random sets; see Theorem 3.2.
Appendix C Computational complexity
The bulk of ShapeDiscover’s computation time is spent on optimization, that is, on executing the $\mathrm{GradientDescent}$ method in Algorithm 3, so let us analyze its computational complexity.
Let $X$ be the input point cloud. The neighborhood graph $G$ built on $X$ has size $|G|=O(\texttt{n\_neigh}\cdot|X|)$. From the loss estimation formulas (Section 4.1), we see that $\widehat{\mathsf{M}}$, $\widehat{\mathsf{G}}$, and $\widehat{\mathsf{R}}$ can be computed in $O(\texttt{n\_cov}\cdot|G|)$ time. A bit more interestingly, the loss $\widehat{\mathsf{T}}_{0}$, which involves the zero-dimensional persistence of a filtered graph, can be computed in $O(\texttt{n\_cov}\cdot|G|\cdot\log|G|)$ time, by sorting edges and using the elder rule (Edelsbrunner & Harer, 2010); see, e.g., Section 5.3.3 in (Rolle & Scoccola, 2024). The overall time complexity is thus:
$O\left(\,\texttt{n\_epoch}\cdot\texttt{n\_cov}\cdot\texttt{n\_neigh}\cdot|X|\cdot\log|X|\,\right).$
The space complexity depends on maintaining the neighborhood graph $G$ and the parameters of the model implementing the fuzzy cover. In the case of the model simply being an $|X|$-by-n_cov real matrix, as in Section 5, the space complexity is $O\left((\texttt{n\_cov}+\texttt{n\_neigh})\cdot|X|\right)$.
Appendix D The filtered simplicial complex associated to a fuzzy cover